OpenStudy (anonymous):

sole the system of equations x+y=5 3x+z=2 4y-z=8

OpenStudy (anonymous):

does anyone know how to do this?

hero (hero):

Eliminate the z variable to create a system of two equations.

OpenStudy (anonymous):

ok so then what?

hero (hero):

If you eliminated z, you should have: x + y = 5 3x + 4y = 10 Now you have a system of two equations and two variables. You should be able to solve that.

OpenStudy (anonymous):

so would that leave z to be 0 then?

hero (hero):

Not necessarily. Solve the system of two equations first. You'll have values for x and y. From there, you'll be able to figure out the value of z

OpenStudy (anonymous):

x=3 and y =2?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

im completely stumped

hero (hero):

I think you are guessing \[3(3) + 4(2) \ne 10\]

hero (hero):

If you cannot solve a system of two equations in two variables, then you should probably learn how to do that before attempting to solve systems of three variables.

OpenStudy (anonymous):

that is why i come here for someone to explain how to solve a system of 2...

OpenStudy (anonymous):

anyway of explaining it?

OpenStudy (anonymous):

3x+4y=10 would i subtract 3 from both sides to find x?

hero (hero):

No, that's not what you do, sorry bro.

OpenStudy (anonymous):

then how do you do it?

OpenStudy (plainntall):

Is this linear algebra?

OpenStudy (anonymous):

no its pre cal

OpenStudy (plainntall):

Okay then solve for one variable in terms of another. x+y=5 3x+z=2 4y-z=8 solve for y in terms of x for x+y=5 then solve for z in terms of x for 3x+z=2 then plug those values into 4y-z=8

OpenStudy (plainntall):

Do you remember to solve equations? You add or multiply both sides of the equation to get only the variable you are solving for on one side

OpenStudy (plainntall):

I'll do the first one x+y=5 x-x+y=5-x ; subtract x from both sides y=5-x ; now you have y in terms of x now solve for z in terms of x for 3x+z=2